OpenStudy (anonymous):

one rectangle has a length of (x +2 )cm and a breadth of 2cm, and another, of equal area, has a length of 5cm and a breadth of (x-3) 1 write down an equation in x and solve it 2 what is the area of each rectangle

OpenStudy (michele_laino):

please you have to solve this equation: \[(x+2)*2=5*(x-3)\]

OpenStudy (anonymous):

help?

OpenStudy (michele_laino):

I'm going to simpify the above equation: \[2x+4=5x-15\] , ading to both sides 15 we have: \[2x+19=5x\] now I add to both sides of last equation -2x, so: \[19=3x\] finallyividing by 3, the last equation, I get: \[x=\frac{ 19 }{ 3 }\]

OpenStudy (anonymous):

what about part 2

OpenStudy (michele_laino):

next step is to insert the value of x, namely 19/3 into the left side, for example, of the first equation, so, we get: \[(x+2)*2=(\frac{ 19 }{ 3 }+2)*2=\frac{ 50 }{ 3 }\] so your area is 50/3 cm^2

OpenStudy (anonymous):

why do we add 2 to 19/3

OpenStudy (michele_laino):

because your rectangle has a length equals to x+2, and, solving our equation, we got x=19/3, so the length of first rectangle is 19/3+2=25/3

OpenStudy (anonymous):

and is this equal to the other rectangle (x-3)

OpenStudy (michele_laino):

not precisely, the area is equal for both rectangles, the single dimensions not, for example the breadth of the second rectangle is x-3=19/3-3=10/3, the breadth of the first rectangle is 2 cm instead, similar for the length

OpenStudy (michele_laino):

for example length of the first rectangle is x+2=19/3+2=25/3, whereas length of the second rectangle is 5 cm

OpenStudy (anonymous):

so the area of the rectangle with the length (x+2) is 50/3cm squared the area of the other rectangle is 20/3 cm squared ?

OpenStudy (michele_laino):

I think no, cause boh areas ae equal to 50/3. Area of secon rectangle is: length*breadth= \[5*(x-3)=5*(\frac{ 19 }{ 3 }-3)=5*\frac{ 10 }{ 3 }=\frac{ 50 }{ 3 }\]

OpenStudy (michele_laino):

please, note that from the text of your problem I read that both rectangles have the same area

OpenStudy (anonymous):

yep thank you ever so much for your help michele !!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

no, thank uuuuuuuuuuuuuuuuuuuuuuuuuu, hope u have a good night, bye